# Theoretical and Empirical Insights into the Origins of Degree Bias in Graph Neural Networks

### v1

@@ 3.2.1 Random Walk Graph Filter
We observe that to make $R_{i,c^{\prime}}$ larger, and thus minimize the probability of misclassification, it is sufficient (although not necessary) that the inverse collision probability $\frac{1}{\sum_{l=0}^{L}\alpha_{i}^{(l)}}$ is larger. When $L=1$, $\frac{1}{\sum_{l=0}^{L}\alpha_{i}^{(l)}}=\frac{1}{\frac{1}{{\bm{D}}_{ii}}+1}$, which is larger for high-degree nodes. We find empirically that the inverse collision probability is positively associated with node degree (cf. Figures 3, 14, 15). Furthermore, disparities in the inverse collision probability across nodes with different degrees is reduced by residual connections and increased by self-loops. Intuitively, random walks starting from high-degree nodes diffuse more quickly, maximizing the probability of any two random walks not colliding at the same end node; in this way, a higher inverse collision probability indicates a more diverse and possibly informative $L$-hop neighborhood. This finding supports hypothesis (H1) (cf. Table 2). Additionally, to make $R_{i,c^{\prime}}$ larger, it is sufficient that for all $l\in[L]$, $\beta_{i,c^{\prime}}^{(l)}$ is more negative, e.g., when most nodes in the $l$-hop neighborhood of $i$ are predicted to belong to class $c$. Thus, $\beta_{i,c^{\prime}}^{(l)}$ can be more negative when nodes in the $l$-hop neighborhood of $i$ also are in class $c$ and were part of the training set $S$, leading to them being correctly classified. Thus, this finding supports hypotheses (H4) and (H6).
@@
Furthermore, our proof of Theorem 2 reveals that in expectation, $\overline{\text{RW}}$ produces similar representations for low and high-degree nodes with similar $L$-hop neighborhood homophily levels. However, low-degree nodes (specifically nodes with a lower inverse collision probability) tend to have a higher variance in $\overline{\text{RW}}$’s representation space than high-degree nodes do. We validate these findings empirically in Figure 2 and §F. In Figure 2, we see in the left plot in the RW row (first row) that low-degree test nodes have representations that are similarly centered but more spread out in the first two principal components of all the test representations than high-degree nodes; we confirm that low-degree node representations have a larger variance in the middle plot in the RW row. Thus, regardless of how RW is trained, low-degree nodes have a higher probability of being on the wrong side of RW’s decision boundaries. Indeed, the left plot in the RW row shows that low-degree nodes of a certain class end up closer to nodes of a different class at a higher rate. Notably, this occurs even when RW is relatively shallow (i.e., 3 layers). Thus, this finding supports hypothesis (H5), as well as draws doubt to hypotheses (H3) and (H10). Our results for $\overline{\text{RW}}$ may also hold for ATT when low-degree nodes are generally less attended to since like random walk transition matrices, attention matrices are row-stochastic.
@@
3.2.2 Symmetric Graph Filter
We now consider the special case that $\forall l\in[L],{\bm{P}}^{(l)}={\bm{P}}_{\text{sym}}={\bm{D}}^{-\frac{1}{2}}{\bm{A}}{\bm{D}}^{-\frac{1}{2}}$. We once again simplify ${\cal M}$ by making all $\sigma^{(l)}$ the identity function, getting $\overline{\text{SYM}}$:
${\bm{H}}^{(L)}=\text{softmax}\left({\bm{Z}}^{(L)}\right)=\text{softmax}\left(\sum_{l=0}^{L}{\bm{P}}^{l}_{\text{sym}}{\bm{X}}{\bm{W}}^{(l)}\right).$
@@ 3.2.2 Symmetric Graph Filter
$\widetilde{\beta}_{i,c^{\prime}}^{(l)}=\mathbb{E}_{j\sim{\cal N}^{(l)}(i)}\left[\frac{1}{\sqrt{{\bm{D}}_{jj}}}\mathbb{E}_{{\bm{x}}\sim{\cal D}_{{\bm{Y}}_{j}}}\left[{\bm{x}}^{T}{\bm{w}}^{(l)}_{c^{\prime}-c}\right]\right]$
@@
$\widetilde{\alpha}^{(l)}_{i}=\sum_{j\in{\cal V}}\frac{1}{{\bm{D}}_{jj}}\left[\left({\bm{P}}_{\text{rw}}^{l}\right)_{ij}\right]^{2},$
@@
Assume that $\forall l\in[L],\forall j\in{\cal V},\text{Var}_{{\bm{x}}\sim{\cal D}_{{\bm{Y}}_{j}}}\left[{\bm{x}}^{T}{\bm{w}}^{(l)}_{c^{\prime}-c}\right]\leq M$. Then:
$R_{i,c^{\prime}}$ $\geq\frac{\left(\sum_{l=0}^{L}\widetilde{\beta}_{i,c^{\prime}}^{(l)}\right)^{2}}{M(L+1)\sum_{l=0}^{L}\widetilde{\alpha}^{(l)}_{i}}.$
Once again, we observe that $R_{i,c^{\prime}}$ is larger, and thus the probability of misclassification is minimized, when the inverse (degree-discounted) sum of collision probabilities $\frac{1}{\sum_{l=0}^{L}\widetilde{\alpha}_{i}^{(l)}}$ is larger and for all $l\in[L]$, the (degree-discounted) $l$-hop prediction homogeneity $\widetilde{\beta}_{i,c^{\prime}}^{(l)}$ is more negative. Like for RW, these findings support hypotheses (H1), (H4), and (H6) (cf. Table 2).
Furthermore, our proof of Theorem 3 reveals that in expectation, $\overline{\text{SYM}}$ often produces representations for low-degree nodes that lie closer to $\overline{\text{SYM}}$’s decision boundary than representations of high-degree nodes with similar $L$-hop neighborhood homophily levels. This is because $\overline{\text{SYM}}$ produces node representations that are approximately scaled by the square root of the node’s degree. However, for the same reason, unlike for $\overline{\text{RW}}$, low-degree nodes tend to have a lower variance in $\overline{\text{SYM}}$’s representation space than high-degree nodes do; this corroborates the findings of [11]. We validate this empirically in Figure 2 and §F on the homophilic datasets (i.e., all datasets except chameleon and squirrel). In Figure 2, we see in the left plot in the SYM row (second row) that low-degree test nodes [^1] Particularly low-degree nodes with many high-degree nodes in their $L$-hop neighborhood. have representations that are closer to SYM’s decision boundaries but less spread out in the first two principal components of all the test representations than high-degree nodes; we confirm that low-degree node representations have a smaller or comparable variance in the middle plot in the SYM row. Our experiments on chameleon and squirrel (cf. §F) do not support our theoretical findings because these datasets have a larger variance in $L$-hop neighborhood heterophily levels across nodes.
@@
3.3 Training-Time Degree Bias
@@
3.3.1 Symmetric Graph Filter
We first demonstrate that during each step of training of $\overline{\text{SYM}}$ with gradient descent, the loss of low-degree nodes is adjusted more slowly than high-degree nodes. We consider the setting that, for all $l\in[L]$, at each training step $t$:
${\bm{W}}^{(l)}[t+1]\leftarrow{\bm{W}}^{(l)}[t]-\eta\frac{\partial\ell[t]}{\partial{\bm{W}}^{(l)}[t]}(B[t]),$
where ${\bm{W}}^{(l)}[t]$ is ${\bm{W}}^{(l)}$ at training step $t$, $\eta$ is the learning rate, $\ell[t]$ is the model’s loss at $t$, and $B[t]\subseteq S$ is the batch used at step $t$.
@@ 3.3.1 Symmetric Graph Filter
$\left|\ell[t+1]({\cal M}|i,c)-\ell[t]({\cal M}|i,c)\right|\leq\left\|{\bm{Z}}_{i}^{(L)}[t+1]-{\bm{Z}}_{i}^{(L)}[t]\right\|_{2}$
Now, we move to the main theorem where we bound the change in loss $i$ after an arbitrary training step $t$ (regardless of batching paradigm) in terms of its degree. We denote the residual of the predictions of $\overline{\text{SYM}}$ at step $t$ by $\epsilon[t]={\bm{H}}^{(L)}[t]-\text{onehot}\left({\bm{Y}}[t]\right)$, where ${\bm{H}}^{(L)}[t]$ and ${\bm{Y}}[t]$ are the submatrices formed from the rows of ${\bm{H}}^{(L)}$ and ${\bm{Y}}$, respectively, that correspond to the nodes in $B[t]$. Furthermore, we define $\forall l\in[L],\widetilde{\chi}_{i}^{(l)}\in\mathbb{R}^{|B[t]|}$, where for $m\in B[t],\left(\widetilde{\chi}_{i}^{(l)}[t]\right)_{m}=\sqrt{{\bm{D}}_{mm}}\mathbb{E}_{j\sim{\cal N}^{(l)}(i),k\sim{\cal N}^{(l)}(m)}\left[\frac{1}{\sqrt{{\bm{D}}_{jj}{\bm{D}}_{kk}}}{\bm{X}}_{j}{\bm{X}}_{k}^{T}\right]$. In effect, $\left(\widetilde{\chi}_{i}^{(l)}[t]\right)_{m}$ captures the degree-discounted expected similarity between the raw features of nodes $j$ and $k$ with respect to the $l$-hop random walk distributions of $i\in{\cal V}$ and $m\in B[t]$.
@@
$\left|\ell[t+1](\overline{\text{SYM}}|i,c)-\ell[t](\overline{\text{SYM}}|i,c)\right|$ $\leq\sqrt{{\bm{D}}_{ii}}\cdot\sqrt{2}\eta\left\|\epsilon[t]\right\|_{F}\sum_{l=0}^{L}\left\|\widetilde{\chi}_{i}^{(l)}[t]\right\|_{2}.$
@@
As observed, the change (either increase or decrease) in loss for $i$ after an arbitrary training step has a smaller magnitude if $i$ is low-degree. Thus, when $\ell[t+1](\overline{\text{SYM}}|i,c)<\ell[t](\overline{\text{SYM}}|i,c)$ (e.g., if $i\in B[t]$), the loss for $i$ decreases more slowly when $i$ is low-degree. In effect, because the magnitude of SYM node representations is positively associated with node degree while the magnitude of each gradient descent step is the same across nodes, the representations of low-degree nodes experience a smaller change during each step. We additionally notice that the loss for $i$ changes more slowly when the features of nodes in its $L$-hop neighborhood are not similar to the features in the $L$-hop neighborhoods of the nodes in each training batch (i.e., $\sum_{l=0}^{L}\left\|\widetilde{\chi}_{i}^{(l)}[t]\right\|_{2}$ is small). Because the $L$-hop neighborhoods of low-degree nodes tend to be smaller than those of high-degree nodes, their neighborhoods often have less overlap with the neighborhoods of training nodes, which can further constrain the rate at which the loss for $i$ changes.

### v2

@@ 4.1 Random Walk Graph Filter
We observe that to make $R_{i,c^{\prime}}$ larger, and thus minimize the probability of misclassification, it is sufficient (although not necessary) that the inverse collision probability $\frac{1}{\sum_{l=0}^{L}\alpha_{i}^{(l)}}$ is larger. When $L=1$, $\frac{1}{\sum_{l=0}^{L}\alpha_{i}^{(l)}}=\frac{1}{\frac{1}{{\bm{D}}_{ii}}+1}$, which is larger for high-degree nodes. We find empirically that the inverse collision probability is positively associated with node degree (see Figures 3, 13, 14). (We elaborate on connections between the inverse collision probability and node degree in §K.) Furthermore, disparities in the inverse collision probability across nodes with different degrees is reduced by residual connections and increased by self-loops. Intuitively, random walks starting from high-degree nodes diffuse more quickly, maximizing the probability of any two random walks not colliding at the same end node; in this way, a higher inverse collision probability indicates a more diverse and possibly informative $L$-hop neighborhood. This finding supports hypothesis (H1) (see Table 2).
Additionally, to make $R_{i,c^{\prime}}$ larger, it is sufficient that for all $l\in\mathbb{N}_{\leq L}$, $\beta_{i,c^{\prime}}^{(l)}$ is more negative, e.g., when most nodes in the $l$-hop neighborhood of $i$ are predicted to belong to class $c$. Thus, $\beta_{i,c^{\prime}}^{(l)}$ can be more negative when nodes in the $l$-hop neighborhood of $i$ also are in class $c$ (i.e., node $i$ has high local homophily) and were part of the training set $S$, leading to them being correctly classified. This finding supports hypotheses (H4) and (H6). Notably, we cannot make $\sum_{l=0}^{L}\beta_{i,c^{\prime}}^{l)}$ more positive to increase $R_{i,c^{\prime}}$; this would violate the assumption of Theorem 2 that the model generalizes in expectation, which is necessary to make a mathematically rigorous statement about degree bias via tail bounds. Intuitively, it also would not make sense that RW and SYM reduce the misclassification error for a node by predicting its neighbors to be of a different class, since message passing smooths the representations of adjacent nodes. Moreover, distribution shifts in local homophily from train to test time can reduce test-time prediction performance, bringing $\beta_{i,c^{\prime}}^{(l)}$ closer to 0; this can increase $R_{i,c^{\prime}}$, thereby not inducing as much degree bias at the expense of overall test performance.
@@
Furthermore, our proof of Theorem 2 (Eqn. 21) reveals that in expectation, the linearized model $\overline{\text{RW}}$ produces similar representations for low and high-degree nodes with similar $L$-hop neighborhood homophily levels. However, low-degree nodes (specifically nodes with a lower inverse collision probability) tend to have a higher variance in $\overline{\text{RW}}$’s representation space than high-degree nodes do (Eqn. 24). This entails that factors beyond homophily (e.g., diversity of neighbors) induce degree bias. We validate these findings empirically in Figure 2 and §F. In Figure 2, we see in the left plot in the RW row (first row) that low-degree test nodes have representations that are similarly centered but more spread out in the first two principal components of all the test representations than high-degree nodes; we confirm that low-degree node representations have a larger variance in the middle plot in the RW row. Thus, regardless of how RW is trained, low-degree nodes have a higher probability of being on the wrong side of RW’s decision boundaries. Indeed, the left plot in the RW row shows that low-degree nodes of a certain class end up closer to nodes of a different class at a higher rate. Notably, this occurs even when RW is relatively shallow (i.e., 3 layers). Thus, this finding supports hypothesis (H5), as well as draws doubt to hypotheses (H3) and (H10). Our results for $\overline{\text{RW}}$ may also hold for ATT when low-degree nodes are generally less attended to since like random walk transition matrices, attention matrices are row-stochastic.
@@
4.2 Symmetric Graph Filter
We now consider the special case that $\forall l\in\mathbb{N}_{\leq L},{\bm{P}}^{(l)}={\bm{P}}_{\text{sym}}={\bm{D}}^{-\frac{1}{2}}{\bm{A}}{\bm{D}}^{-\frac{1}{2}}$. We once again simplify ${\cal M}$ by making all $\sigma^{(l)}$ the identity function, getting $\overline{\text{SYM}}$:
${\bm{H}}^{(L)}=\text{softmax}\left({\bm{Z}}^{(L)}\right)=\text{softmax}\left(\sum_{l=0}^{L}{\bm{P}}^{l}_{\text{sym}}{\bm{X}}{\bm{W}}^{(l)}\right).$ (7)
@@ 4.2 Symmetric Graph Filter
$\widetilde{\beta}_{i,c^{\prime}}^{(l)}=\mathbb{E}_{j\sim{\cal N}^{(l)}(i)}\left[\frac{1}{\sqrt{{\bm{D}}_{jj}}}\mathbb{E}_{{\bm{x}}\sim{\cal D}_{{\bm{Y}}_{j}}}\left[{\bm{x}}^{T}{\bm{w}}^{(l)}_{c^{\prime}-c}\right]\right]$ (8)
@@
$\widetilde{\alpha}^{(l)}_{i}=\sum_{j\in{\cal V}}\frac{1}{{\bm{D}}_{jj}}\left[\left({\bm{P}}_{\text{rw}}^{l}\right)_{ij}\right]^{2},$ (9)
@@
Assume that $\forall l\in\mathbb{N}_{\leq L},\forall j\in{\cal V},\text{Var}_{{\bm{x}}\sim{\cal D}_{{\bm{Y}}_{j}}}\left[{\bm{x}}^{T}{\bm{w}}^{(l)}_{c^{\prime}-c}\right]\leq M$. Then:
$R_{i,c^{\prime}}$ $\geq\frac{\left(\sum_{l=0}^{L}\widetilde{\beta}_{i,c^{\prime}}^{(l)}\right)^{2}}{M(L+1)\sum_{l=0}^{L}\widetilde{\alpha}^{(l)}_{i}}.$ (10)
Once again, we observe that $R_{i,c^{\prime}}$ is larger, and thus the probability of misclassification is minimized, when the inverse (degree-discounted) sum of collision probabilities $\frac{1}{\sum_{l=0}^{L}\widetilde{\alpha}_{i}^{(l)}}$ is larger and for all $l\in\mathbb{N}_{\leq L}$, the (degree-discounted) $l$-hop prediction homogeneity $\widetilde{\beta}_{i,c^{\prime}}^{(l)}$ is more negative. Like for RW, these findings support hypotheses (H1), (H4), and (H6) (see Table 2).
Furthermore, our proof of Theorem 3 (Eqn. 33) reveals that in expectation, $\overline{\text{SYM}}$ often produces representations for low-degree nodes that lie closer to $\overline{\text{SYM}}$’s decision boundary than representations of high-degree nodes with similar $L$-hop neighborhood homophily levels. This is because $\overline{\text{SYM}}$ produces node representations that are approximately scaled by the square root of the node’s degree. However, for the same reason, unlike for $\overline{\text{RW}}$, low-degree nodes tend to have a lower variance in $\overline{\text{SYM}}$’s representation space than high-degree nodes do (Eqn. 36); this corroborates the findings of [12]. We validate this empirically in Figure 2 and §F on the homophilic datasets (i.e., all datasets except chameleon and squirrel). In Figure 2, we see in the left plot in the SYM row (second row) that low-degree test nodes (particularly low-degree nodes with many high-degree nodes in their $L$-hop neighborhood) have representations that are closer to SYM’s decision boundaries but less spread out in the first two principal components of all the test representations than high-degree nodes; we confirm that low-degree node representations have a smaller or comparable variance in the middle plot in the SYM row. We emphasize that while SYM representations of high-degree nodes have a higher variance, this itself is not the cause of degree bias; since the standard deviation and expectation of SYM node representations are approximately scaled by the same factor, by Theorem 1, the variance of SYM representations of high-degree nodes does not enlarge $R_{i,c^{\prime}}$ noticeably more than in the RW case.
Notably, our theoretical findings do extend to heterophilic graphs. In particular, high-degree nodes in heterophilic networks (e.g., chameleon and squirrel) do not have higher negative $L$-hop prediction homogeneity levels due to higher local heterophily (see §F), and hence we do not necessarily observe better test performance for them (see Figure 5). None of our theoretical analysis assumes homophilic networks.
@@
5 Training-Time Degree Bias
@@
We first demonstrate that during each step of training of $\overline{\text{SYM}}$ with gradient descent, the loss of low-degree nodes is adjusted more slowly than high-degree nodes. We consider the setting that, for all $l\in\mathbb{N}_{\leq L}$, at each training step $t$:
${\bm{W}}^{(l)}[t+1]\leftarrow{\bm{W}}^{(l)}[t]-\eta\frac{\partial\ell[t]}{\partial{\bm{W}}^{(l)}[t]}(B[t]),$ (11)
where ${\bm{W}}^{(l)}[t]$ is ${\bm{W}}^{(l)}$ at training step $t$, $\eta$ is the learning rate, $\ell[t]$ is the model’s loss at $t$, and $B[t]\subseteq S$ (where $S\subseteq{\cal V}$ is the labeled subset of nodes) is the batch used at step $t$.
@@
$\left|\ell[t+1]({\cal M}|i,c)-\ell[t]({\cal M}|i,c)\right|\leq\left\|{\bm{Z}}_{i}^{(L)}[t+1]-{\bm{Z}}_{i}^{(L)}[t]\right\|_{2}$ (12)
Now, we move to the main theorem where we bound the change in loss $i$ after an arbitrary training step $t$ (regardless of batching paradigm) in terms of its degree. We denote the residual of the predictions of $\overline{\text{SYM}}$ at step $t$ by $\epsilon[t]={\bm{H}}^{(L)}[t]-\text{onehot}\left({\bm{Y}}[t]\right)$, where ${\bm{H}}^{(L)}[t]$ and $\text{onehot}\left({\bm{Y}}[t]\right)$ are the submatrices formed from the rows of ${\bm{H}}^{(L)}$ and $\text{onehot}\left({\bm{Y}}\right)$, respectively, that correspond to the nodes in $B[t]$. Furthermore, we denote $\forall l\in\mathbb{N}_{\leq L}$, the expected similarity of the neighborhoods of $i$ and $B[t]$ by $\widetilde{\chi}_{i}^{(l)}\in\mathbb{R}^{|B[t]|}$, where for $m\in B[t],\left(\widetilde{\chi}_{i}^{(l)}[t]\right)_{m}=\sqrt{{\bm{D}}_{mm}}\mathbb{E}_{j\sim{\cal N}^{(l)}(i),k\sim{\cal N}^{(l)}(m)}\left[\frac{1}{\sqrt{{\bm{D}}_{jj}{\bm{D}}_{kk}}}{\bm{X}}_{j}{\bm{X}}_{k}^{T}\right]$. Specifically, $\left(\widetilde{\chi}_{i}^{(l)}[t]\right)_{m}$ captures the degree-discounted expected similarity between the raw features of nodes $j$ and $k$ with respect to the $l$-hop random walk distributions of $i\in{\cal V}$ and $m\in B[t]$. Notably, our matrix is pre-feature aggregation (e.g., unlike [37]).
@@
$\left|\ell[t+1](\overline{\text{SYM}}|i,c)-\ell[t](\overline{\text{SYM}}|i,c)\right|$ $\leq\sqrt{{\bm{D}}_{ii}}\cdot\sqrt{2}\eta\left\|\epsilon[t]\right\|_{F}\sum_{l=0}^{L}\left\|\widetilde{\chi}_{i}^{(l)}[t]\right\|_{2}.$ (13)
@@
As observed, the change (either increase or decrease) in loss for $i$ after an arbitrary training step has a smaller magnitude if $i$ is low-degree. Thus, when $\ell[t+1](\overline{\text{SYM}}|i,c)<\ell[t](\overline{\text{SYM}}|i,c)$ (e.g., if $i\in B[t]$), the loss for $i$ decreases more slowly when $i$ is low-degree. In effect, because the magnitude of SYM node representations is positively associated with node degree while the magnitude of each gradient descent step is the same across nodes, the representations of low-degree nodes experience a smaller change during each step. We additionally notice that the loss for $i$ changes more slowly when the features of nodes in its $L$-hop neighborhood are not similar to the features in the $L$-hop neighborhoods of the nodes in each training batch (i.e., $\sum_{l=0}^{L}\left\|\widetilde{\chi}_{i}^{(l)}[t]\right\|_{2}$ is small). Because the $L$-hop neighborhoods of low-degree nodes tend to be smaller than those of high-degree nodes, their neighborhoods often have less overlap with the neighborhoods of training nodes, which can further constrain the rate at which the loss for $i$ changes. Notably, while node degree highly affects the rate of learning, differences in $\widetilde{\chi}$ across nodes due to factors other than degree are also influential.